OpenStudy (anonymous):

Solve the equation for x: 2x – 14y = 2

jhonyy9 (jhonyy9):

2x -14y=2 2(x-7y)=2 divide both sides by 2 x-7y=1 x=7y+1

OpenStudy (anonymous):

thanks